Pharoah Company, organized in 2019, has set up a single account for all intangible assets. The following summary discloses the debit entries that have been recorded during 2020.

1/2/20 Purchased patent (7-year life)

$304,500

4/1/20 Purchase goodwill (indefinite life)

345,000

7/1/20 Purchased franchise with 10-year life; expiration date 7/1/30

425,000

8/1/20 Payment of copyright (5-year life)

150,000

9/1/20 Research and development costs

215,000

$1,439,500

Prepare the necessary entry to clear the Intangible Assets account and to set up separate accounts for distinct types of intangibles. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Make the entry as of December 31, 2020, recording any necessary amortization. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Reflect all balances accurately as of December 31, 2020.

Use straight-line amortization

select an intangible asset

$enter a dollar amount

select an intangible asset

$enter a dollar amount

select an intangible asset

$enter a dollar amount

select an intangible asset

$enter a dollar amount
Answer #1
1
Account Titles and Explanation Debit Credit
Patents 304500
Goodwill 345000
Franchises 425000
Copyrights 150000
Research and Development Expense 215000
         Intangible Assets 1439500
2
Account Titles and Explanation Debit Credit
Amortization Expense 77250
      Patents 43500 =304500/7
      Franchise 21250 =425000/10*6/12
      Copyright 12500 =150000/5*5/12
3
Use straight-line amortization
Patents 261000 =304500-43500
Goodwill 345000
Franchise 403750 =425000-21250
Copyright 137500 =150000-12500